OpenStudy (anonymous):

Converting angular accel. units: When going from rad/min^2 to rad/s^2, would I simply divide by 60, or will I need to account for the 'squared' somehow?

OpenStudy (anonymous):

\[\frac{ rad }{ \min ^{2} }=\frac{ rad }{ \min \times \min }\] \[\frac{ rad }{ 60s \times 60s }=\frac{ rad }{3600 s ^{2} }\] \[\frac{ 1 }{ 3600}\times \frac{ rad }{ s ^{2} }\] you actually divide by 3600